A bank loaned out $14,000, part of it at the rate of 9% per year and the rest at 17% per year. If the interest received in one year is $2000, how much was loaned at 9%?
Answer by stanbon(75887)

A bank loaned out $14,000, part of it at the rate of 9% per year and the rest at 17% per year. If the interest received in one year is $2000, how much was loaned at 9%?
-----
Equations:
n + s = 14000
9n + 17s = 200000
-------------------------
Modify for elimination:
17n + 17s = 17*14000
9n + 17s = 200000
-----------------------
Subtract and solve for "n"::
8n = 38000
n = $4750 (amt invested at 9%)
------------
Cheers,
Stan H.
-------------------